From GlueXWiki
Jump to: navigation, search

FDC Weekly Meeting

Date: December 4, 2008

Participants: Daniel, Simon, Brian, Bill, Mark S., Micah, Herun, Mark I., Chuck, Roger, Beni, David, Fernando, Jim

Next Meeting: December 11, 2008 @ 1:30 p.m.

Full-Scale Prototype Updates

 - Procurements:
    * Rigid-flex assemblies:
       * Fernando and Roger have not yet converged on review/approval of
         the design drawings. The plan (like it was at our last meeting)
         to move forward with a PR for a small number of boards in the
         next couple of days. The lead time is about 3 weeks, so the 
         boards will not be in our hands this calendar year. DSC stressed
         the importance of completing this order and getting the main
         order out the door as soon as we can.
       * Roger has passed on the board stuffing details to Kim and she
         has communicated with Greg Arnold's group.
    * Wire QA studies:
       * The QA checks for the sense wires are complete.
       * Brian has prepared a slide for the SEM studies of the field wire
         and given it to the SEM folks. They will complete their work in
         about a week. The mechanical load studies of the field wires were
         done and the wire passed (barely).
       * Brian will prepare a final QA report with photos and test results
         for all wires in the next week.
       * Brian checked with the purchasing agent responsible for the
         wire order to find out what QA qualifications were part of this
         order if any. Apparently they were not part of the order. It is
         not clear what happened. Was this a miscommunication or a did
         somebody drop the ball?
    * G10 ring order status:
      * No new news. Expected delivery date is roughly 1 week.
    * 110IG Rohacell order status:
      * No new news. Expected delivery date is Dec. 17.
 - Mark S. has taken responsibility for watching over the FDC component
   stores and keeping the database up to date. 
 - Mark S. is also responsible for upkeep of the list of procurements for
   the full-scale prototypes.
      * Location: M Drive/halld/Electronics, or for more sophisticated folks, 
 - Bill is working to put the local HV connectors selected by Fernando into 
   the 3D CAD model and to detail how they will be attached (as they are 
   quite massive).  Bill will get back to this work when he has a chance.

Wire Frame Construction

 - Bill has completed a draft of a specification/tolerance document
   for the wire frames and we discussed it. Presently the final tolerance
   is given as a linear sum of the different sources of uncertainty. It is
   more likely relevant to quote the quadrature sum.
 - Mark S. and Brian have completed assembly of the 4 PCB annuli. This
   included all touch up work and cleaning. The board rings were placed in
   the plexiglas storage box for safety.
 - Bill has worked with the JLab machine shop to do some tests of the 
   system to do vacuum bagging on the G10/foam laminate. While things did
   not work fully has Bill had hoped, he declared that we are ready for
   preparation of the G10/foam laminates once the orders arrive. Bill
   has already provided the drawing for the foam cutting template to the
   machine shop (remember that we will be piecing together our foam cores
   from several arcs).
 - Bill has communicated with the machine shop about the machining
   requirements of our G10/foam laminates. DSC has a remaining question on
   whether any machining is needed on the wire frames after assembly. We
   should be able to get this work through the machine shop relatively
 - Mark S. is working to modify the wire frame construction procedures 
   document including the photographs that he took. A new draft has been
   passed around to Bill and Brian for feedback. A draft will be given to
   DSC shortly for review.
 - DSC talked to Keith at IUCF with regard to the unofficial schedule
   for winding.  It looks like we should be able to get the wire frames
   to IUCF by the middle of January if our current schedule holds up.
 - Kim has provided a draft of a procedures document for the Phase 1 and
   Phase 2 wire frame PCB component attachment. She has merged this
   document with Fernando's cleaning and handling document.  The document
   is being finalized with some additional language on soldering iron
   temperature vs. time profiles.  It should be completed before the
   start of the Phase 2 component soldering.

Cathode Planning

 - We reviewed the cathode R&D timeline. Folks are reminded to pass on
   progress updates to Chuck every Wednesday.  The current schedule is
   being kept on the mdrive at:
     * Location: M:\halld\Individual-Schedules\FDC or for more sophisticated 
       folks, /u/group/halld/Individual-Schedules/FDC
 - Bill has completed a draft of a specification/tolerance document
   for the cathode frames and we discussed it. Presently the final tolerance
   is given as a linear sum of the different sources of uncertainty. It is
   more likely relevant to quote the quadrature sum.
 - Bill has completed the design of the cathode edge cutting system. He is
   finalizing the fabrication drawings and then the procurement will begin. 
   Some of the parts actually have already been ordered and delivered to
   JLab. After assembly of the system, we will work through a full test
   plan before cutting any of our "real" cathodes.
 - Cathode flatness measuring system:
     * It was verified that the problems seen in the scan data presented
       last week by Micah were due to rail misalignment problems. The rail
       was reset and the granite table surface rescanned.  It seems that a
       rail misalignment still remains, although at about half the level
       it was at before.
     * Bill arranged for the Survey Group to survey the flatness of the
       granite table (to separate problems with the table surface from
       rail misalignment issues). The work will be completed by the end
       of the day today. The survey data will then be passed along to Bill.
     * David showed us comparisons of table scans before and after resetting
       the rails.  See the FDC wiki for details.
     * Micah and David are working on documentation for the system to 
       provide full instructions for operation and data analysis. Mark S.
       will take over as the system expert and Micah will transfer knowledge
       to him.
     * Micah and Mark S. will be responsible for completing the measurements
       on the high tension cathode as a function of rotation angle.  This
       work should be completed in the next week.
 - Still waiting on an updated version of the cathode construction document.
   Bill has this now and will get back to it soon.
 - Brian and Mark S. have made up a new cathode prototype using the 5 micron
   Cu dummy boards. This prototype was built for several reasons:
     * To remind folks of how to make the cathodes seeing that one has not
       been built for some time. 
     * To test out the new glue strips. Two-inch wide Kapton strips painted
       with high strength hysol adhesive using the foam rollers. The technique
       worked beautifully and the epoxy stayed where it was supposed to.
     * To set up a system to measure the cathode tension in the tensioning
       system. Brian rigged up a bar to support the laser head over the
       center of the cathode. With the help of Eugene that measured the
       tension in situ. Eugene has completed his tension measurements and
       has placed his results on his web site. 
       The measurements demonstrate a systematic dependence of the result on 
       the radius of about 4%. The measurements with different weights are 
       consistent within 1%, apart from one point, which must be an error of 
       the particular measurement (the laser moved?). The foil elongation 
       perpendicular to the strips was measured to be about 0.33% for the
       1-mil Kapton with 5-micron Cu at about 1300 N/m. 
     * Brian has attached the cathode to a G10 ring and will leave it for the 
       next several weeks to make sure that there are no problems with creep.
     * Bill is still working on the design for the system to hold the laser
       head and allow movement for measurements at several locations. It is
       presently not clear if this is a necessity.  DSC willl  talk to Eugene.
     * Brian has been asked to come up with a detailed list of remaining
       prototypes that need to be made before we can consider making up a
       cathode from our "full-scale prototype lot".
 - We will use the sample board from Allflex (2-micron Cu) for performing
   some handling studies. We will then use it to measure the stretching of
   the foil at the nominal cathode tension to compre the properties of
   Espinex-m to Kapton.
 - Mark S. made up several prototypes of a scheme to connect the cathode
   boards to the external ground. Folks should look these over from the
   standpoint of mechanical and electrical connections and give feedback.

Spacer Design

 - Bill has declared that the spacer ring will have two o-ring grooves.
   This is the best solution (and most viable). He will move to have
   the machine shop prepare the polyethylene and poly propylene rings.
   These will be complete spacers, and thus, will include the o-ring
   grooves and the gas plumbing.

Stack Assembly Procedures

 - DSC, Brian, and Bill will work to finalize the stack assembly construction 
   document that has been prepared.  This document should be in place before 
   we get too far in the construction process.
 - Bill has found some plastic threaded rods that could be used to 
   compress the stack.  We will be receiving some samples shortly and
   will carry out some loading tests.

Cooling System Tests

 - Fernando is in the process of writing up the results from his cooling 
   system studies.  Stay tuned for the GlueX note.
 - Bill has done some ANSYS modeling to better study the connector that 
   goes from the daughter board card to the cooling loop.  These will be
   added to Fernando's note.

Gas System Design

 - Slava has worked on a preliminary layout of the gas handling system 
   for the CDC and FDC.  He provided two options for the FDC system design 
   that were essentially a basic version and a fancier version.  Slava's 
   design work was based on the original concept that Brian came up with 
   long ago.  We will continue the design work as we have time.


 - Roger needs to prepare a document for QA/stuffing/cleaning for the
   cathode boards and a similar document for the cathode daughter
   boards and ground boards.

Small-Scale Prototype

 - We will be moving the FDC cosmic ray test stand from the Test Lab back
   to the EEL building. This move should be completed in the next 6 months
   or so.

Work List

 - The FDC short-term work list has been posted on the FDC web site.  
   This is continually being updated and DSC welcomes any feedback or comments 
   from the group.

Minutes prepared by Daniel. Send any comments or corrections along.